### v1

Abstract
We introduce OneCAT, a unified multimodal model that seamlessly integrates understanding, generation, and editing within a novel, pure decoder-only transformer architecture. Our framework uniquely eliminates the need for external components such as Vision Transformers (ViT) or vision tokenizer during inference, leading to significant efficiency gains, especially for high-resolution inputs. This is achieved through a modality-specific Mixture-of-Experts (MoE) structure trained with a single autoregressive (AR) objective, which also natively supports dynamic resolutions. Furthermore, we pioneer a multi-scale visual autoregressive mechanism within the Large Language Model (LLM) that drastically reduces decoding steps compared to diffusion-based methods while maintaining state-of-the-art performance. Our findings demonstrate the powerful potential of pure autoregressive modeling as a sufficient and elegant foundation for unified multimodal intelligence. As a result, OneCAT sets a new performance standard, outperforming existing open-source unified multimodal models across benchmarks for multimodal generation, editing, and understanding.
1 Introduction
In the past few years, modular approaches, utilizing separate architectures for understanding [2, 65, 3, 13, 14, 12], generation [40, 36, 59, 24, 34, 20], and editing tasks [85, 5, 32, 87, 46, 86], dominated multi-modal frameworks. While producing capable systems, it inherently creates complex, multi-stage pipelines. Such designs often suffer from architectural bottlenecks that limit deep, early-stage fusion of cross-modal information and introduce significant inference latency, presenting a major barrier to both efficiency and performance. In response to these limitations, the field has seen a rapid convergence towards unified multimodal modeling, aiming to integrate these disparate capabilities within a single, end-to-end architecture [67, 62, 75, 76, 11, 16]. Despite the trend towards unification, many current models remain tethered to the modular paradigm. We contend that a true paradigm shift—necessary to unlock the full potential of unified systems—demands a move towards a more fundamental, first-principles architecture that eschews external components. We therefore propose that a pure decoder-only transformer, trained under a unified objective, provides not just a sufficient, but a more elegant and potent foundation for the next generation of general-purpose multimodal intelligence.
Figure 1: Showcase of the text-to-image generation abilities of the OneCAT model.
Figure 2: Showcase of the image editing abilities of the OneCAT model, including general image editing tasks such as object removal, background adjustment, color adjustment, subject replacement, and style transfer; as well as perceptual tasks including depth estimation, pose estimation, object segmentation, and Canny edge detection.
To realize this vision, we re-evaluate the core architectural tenets of multimodal systems. We introduce an encoder-free framework where raw visual inputs are directly tokenized into patch embeddings and processed alongside text tokens within a single decoder stack. The critical innovation is a modality-specific Mixture-of-Experts (MoE) layer, which dynamically routes vision and text tokens to specialized experts, enabling deep, early-stage feature fusion in an efficient inference manner without the need for exquisite encoders. For generative tasks, we pioneer a multi-scale autoregressive mechanism [24, 63] inside the LLM, augmented with the proposed scale-aware adapter modules, to predict image tokens from low to high resolution. This design not only circumvents the high latency of iterative diffusion models but also allows the model to learn a coarse-to-fine generative process, significantly enhancing both speed and output quality.
Regarding data protocol, we leverage mixed training strategies that combine large-scale, web-scraped image-text pairs with curated, instruction-following datasets. Our pre-training corpus is constructed to expose the model to a diverse range of data formats, including image-text documents, visual question-answering pairs, and generative prompts for both text-to-image synthesis and image editing. This heterogeneous data mixture is crucial for training a truly unified model, as it forces the shared decoder to develop a generalized representation that can seamlessly switch between comprehension, generation, and editing tasks. We argue that this unified data diet, coupled with our bottleneck-free architecture, is key to fostering the synergistic emergence of complex multimodal capabilities.
Building upon these principles, we present Only DeCoder Auto-regressive Transformer (OneCAT), an open-source unified multimodal model. Comprehensive evaluations demonstrate that OneCAT sets a new state-of-the-art for pure decoder-only unified models. More importantly, the encoder-free design provides a significant inference speedup, particularly for high-resolution inputs. By demonstrating the viability and superiority of a pure decoder architecture, OneCAT offers a more first-principles-aligned paradigm for multimodal modeling. It facilitates earlier cross-modal fusion through its unified MoE structure and enhances semantic consistency, providing valuable insights and a powerful new baseline for the development of next-generation unified multimodal systems. For examples of our model’s impressive image generation capabilities, please refer to Fig. 1. Its advanced image editing functionalities are further showcased in Fig. 2.
2 Related Work
2.1 Compositional MLLMs
The field of Multimodal Large Language Models (MLLMs) has rapidly evolved, converging on a dominant compositional architecture. This paradigm connects a pre-trained vision encoder, such as CLIP [56], SigLIP [84], or the more recent InternViT [14], to a powerful Large Language Model (LLM) through a trainable connector. Pioneering works [1, 37] propose sophisticated connector designs. For example, Flamingo [1] introduces gated cross-attention layers to inject visual information into an LLM, while BLIP-2 [37] develops the Q-Former to bridge the modality gap between an image encoder and an LLM. A significant shift occurs with LLaVA [45], which simplifies the connector to a lightweight Multi-Layer Perceptron (MLP) projection layer. This effective design become a foundational blueprint for numerous subsequent MLLMs. For example, recent state-of-the-art models like the InternVL series [14, 12, 13, 91, 66] and the Qwen-VL series [65, 3, 2] adopt this same core principle and achieve superior performance by leveraging larger-scale training data and more powerful vision and language foundation models. However, this successful compositional design has inherent drawbacks. The separate nature of the vision and language components complicates the end-to-end optimization process and introduces two critical bottlenecks. First, the sequential nature of the architecture, where the vision encoder must fully process an image before the LLM can begin its generation, leads to high inference latency. This initial step is often referred to as the “prefilling” stage. Second, the connector itself acts as an information bottleneck. In this so-called late fusion pipeline, complex visual information is compressed into a compact representation for the LLM, inevitably causing a loss of fine-grained visual detail. These fundamental limitations are now motivating a shift in the field towards more deeply integrated, such as decoder-only models, that aim to overcome these challenges.
2.2 Decoder-only MLLMs
Decoder-only MLLMs, also known as monolithic MLLMs, have recently emerged as a minimalist yet powerful alternative to the conventional compositional architecture. This paradigm aims to achieve greater efficiency and a more direct early fusion of modalities by eliminating the separate vision encoder. For example, Fuyu-8B [4] processes vision patches by feeding them through a simple linear patch embedding layer directly into the LLM backbone, which markedly reduce inference latency. Inspired by this success, subsequent works [18, 49, 19] further advance decode-only MLLMs by targeting their training processes and architectures. EvE [18] introduces a novel training objective that guides the model’s visual learning by aligning the last layer of LLM’s hidden states of image patch tokens with semantic features from a pre-trained vision encoder. Differently, Mono-InternVL [49] and EvEv2.0 [19] adopt a Mixture-of-Experts (MoE) framework, introducing a dedicated visual expert to handle visual-specific features more effectively. HoLVE [61] prepends a causal transformer to the LLM to explicitly convert both visual and textual inputs into a shared space. Despite these promising advancements, the overall training efficiency of these models remains a significant challenge. More importantly, the potential for the decoder-only architecture to create unified models that can seamlessly integrate multimodal understanding, generation, and even image editing capabilities remains a largely unexplored research avenue.
2.3 Unified Visual Understanding and Generation
Building on the success of MLLMs, the convergence of visual understanding and generation into a unified framework now represents a key research frontier. Pioneering unified MLLMs such as Chameleon [62],Transfusion [90], emu3 [67] and show-o [75] utilize visual tokenizers (e.g., VQ-VAE) to convert images into a sequence of discrete tokens, enabling seamless multimodal understanding and generation within a single model. However, the discretization inevitably results in lossy visual information and weakens in extracting semantic contents. Janus series [69, 11] decouples visual encoding for understanding and generation using two separate encoders, but may compromise performance due to task conflicts in shared LLM parameter space. Metaqueries [53], BLIP3-O [8], Uniworld-V [43] assembles off-the-shelf specialized MLMMs and diffusion models by tuning adapters and learnable query tokens, which sacrifices true architectural unification for modularity. BAGEL [16] and Mogao [41] employ a Mixture-of-Transformers (MoT) architecture, dedicating different components to autoregressive text generation and diffusion-based visual generation. However, while powerful, this hybrid approach inherits the significant inference latency of diffusion models and still requires separate encoders and tokenizers during the inference.
In contrast to these approaches, our OneCAT introduces a pure decoder-only architecture. By integrating modality-specific experts directly within the decoder, OneCAT achieves versatile multimodal capabilities without the need for external vision encoders or tokenizers at inference time, thus resolving the trade-off between architectural purity and inference efficiency.
2.4 Next Scale Prediction for Visual Generation
Autoregressive models based on next-token prediction(NTP) have long faced efficiency challenges in high-resolution image generation due to the quadratic growth of sequence length with image size. Similarly, diffusion models—though widely successful—often suffer from slow iterative sampling. To address these limitations, VAR [63] introduced the next-scale prediction(NSP) paradigm, which encodes images into hierarchical discrete tokens via a multi-scale VAE and generates them autoregressively from low to high resolution, significantly reducing the number of decoding steps. Building upon this, Infinity [24] further enhanced this approach with bit-level prediction and extended tokenizer vocabulary, achieving superior generation quality while maintaining efficient inference. To enable unified understanding and generation, VARGPT [92] stack the transformer from pretrained VAR [63] as a visual decoder atop a LLM. However, since the visual tokens (i.e., the input of the visual decoder) must be decoded token-by-token through the LLM before subsequent next-scale prediction, this approach compromises the inference efficiency that is the key advantage of the NSP.
In contrast, our proposed OneCAT seamlessly unifies next-token prediction for text generation and next-scale prediction for visual generation within a single LLM.
3 OneCAT
Figure 3: Overview inference pipeline of OneCAT. OneCAT is a decoder-only autoregressive unified model that seamlessly supports multimodal understanding, text-to-image generation and image editing. OneCAT implements both next-token prediction for text generation and next-scale prediction for visual generation within a unified LLM backbone. In the left part, the dash squares denote the generated text and visual token during inference. In the right part, PE and $T$ denote patch embedding and text tokenizer, respectively. TD and ID denote text detokenizer and image detokenizer, respectively.
As illustrated in Fig. 3, OneCAT employs a pure decoder-only architecture, eliminating the need for any additional vision encoder or tokenizer during inference. This streamlined design significantly simplifies the model structure and reduces computational overhead. Unlike traditional multimodal models that rely on semantic encoders like ViTs for understanding [3, 16, 41, 11], OneCAT utilizes a lightweight Patch Embedding layer. This layer losslessly converts raw images into continuous visual tokens, enabling efficient multimodal understanding. Crucially, the same Patch Embedding layer also encodes reference images for editing tasks, thereby superseding separate VAE-based tokenizers traditionally used [16, 41] and further enhancing inference efficiency.
At its core, OneCAT integrates a Mixture-of-Experts (MoE) architecture. This MoE comprises three specialized feed-forward network (FFN) experts: one dedicated to processing text tokens for language comprehension, another designed for continuous visual tokens to facilitate multimodal understanding, and a third optimized for generating discrete visual tokens during image synthesis. All QKV and attention layers are shared across modalities and tasks, which promotes significant parameter efficiency and robust cross-modal alignment for instruction-following.
For text generation, OneCAT adheres to the conventional Next-Token Prediction paradigm, leveraging the well-established capabilities of autoregressive language modeling. In parallel, for visual generation, it innovatively employs the Next-Scale Prediction paradigm [63, 24]. This mechanism generates images in a coarse-to-fine, hierarchical manner, progressively predicting visual tokens from the lowest to the highest resolution scale, thereby achieving high-quality visual outputs.
3.1 Architecture
Our OneCAT model is initialized from the pre-trained Qwen2.5 LLM [77], leveraging its strong language modeling capabilities as a foundation. To construct our MoE architecture, we replicate the original FFN layer from each Qwen2.5 transformer block to form three distinct experts: a Text FFN (i.e., Text. FFN), a Visual Understanding FFN (i.e., Und. FFN), and a Visual Generation FFN (i.e., Gen. FFN). OneCAT employs a deterministic hard routing mechanism, where tokens are assigned to a specific expert based on their modality and the task at hand. The core functionality for each task is handled as follows:
Multimodal Understanding.
To process images for understanding tasks, we employ a simple yet effective patch embedding layer that converts raw images into a sequence of continuous visual tokens. This layer consists of a $\times14$ convolution for image patchifying, a $\times2$ pixel unshuffle operation for visual token compression, and a two-layer Multilayer Perceptron (MLP) for projecting the visual features to match the LLM’s hidden state dimension. These continuous visual tokens are exclusively routed to the Und. FFN, while the text tokens for instruction are routed to the Text FFN.
Text-to-Image Generation.
For image generation, we leverage a pre-trained multi-scale VAE model from Infinity [24] to map images between pixel space and latent space. This VAE operates with a downsampling ratio of 16 and a latent channel size of 32, and incorporates a bitwise quantizer [89] to enlarge the vocabulary. During training, the image tokenizer transforms the target images into a sequence of multi-scale discrete visual tokens to serve as ground-truth and input of LLM for teacher-forcing training, which are processed by the Gen. FFN. Critically, during inference, the tokenizer is not required; only the detokenizer is needed to reconstruct the final image from the generated multi-scale visual tokens. The conditional text tokens are also routed to the Text FFN.
Image Editing.
OneCAT seamlessly supports image editing task by conditioning the visual generation process on a reference image and edit-instruction. The reference image is processed through the patch embedding layer, and the resulting continuous visual tokens are routed to the Und. FFN to serve as the visual condition, while the text tokens for instruction are routed to the Text FFN. The patch embedding layer provides a near-lossless representation of the reference image, which allows the LLM’s shallower layers to access low-level features for pixel-level consistency, while the deeper layers extract high-level features for semantical comprehension. Guided by this rich, hierarchical visual context, the model then autoregressively predicts new discrete visual tokens, which are handled by the Gen. FFN. This design enables powerful conditional generation without requiring any architectural modifications, showcasing the versatility of our unified decoder-only design.
3.2 Scale-Aware Adapter for Hierarchical Generation
While the Gen. FFN is capable of processing discrete visual tokens, the tokens produced by the multi-scale VAE are inherently hierarchical. A standard FFN would treat tokens for different scales equally, ignoring this crucial structural information. To address this and enable more granular control over the visual generation process, we introduce the Scale-Aware Adapter (SAA), a novel architectural component integrated with the Gen. FFN.
Our design is motivated by the principle that different scales in the multi-scale VAE govern distinct aspects for image generation. Specifically, tokens from lower scales predominantly encode low-frequency global information, such as color, illumination, and coarse structure. Those from higher scales, conversely, capture high-frequency details including fine textures and intricate patterns. Processing these functionally divergent tokens with a shared Gen. FFN layer limits representational capacity and is thus suboptimal.
The Scale-Aware Adapter (SAA) comprises a set of parallel modules that serve as skip connections over the Gen. FFN. Each module is dedicated to processing tokens from a specific scale of the multi-scale VAE, with the total number of adapters matching the number of VAE scales. During inference, discrete visual tokens are routed to their corresponding scale-specific adapter based on the scale index. To ensure parameter efficiency, each adapter is constructed using a low-rank decomposition (rank rr= 64), inspired by the LoRA [26]. However, unlike LoRA which is typically used for fine-tuning, the SAA modules are trained jointly end-to-end as permanent components of the LLM.
3.3 Multimodal versatile attention mechanism
We leverage a multimodal versatile attention mechanism based on PyTorch FlexAttention [55] to empower OneCAT with the ability to process diverse modalities and tasks in a flexible and adaptive manner. As illustrated in Fig. 4, text tokens $T$ are processed using causal attention, ensuring autoregressive generation; Continuous visual tokens $U$ are processed via full attention, allowing each token to attend to all others in the sequence. Multiscale discrete visual tokens $G_{i}$ (where $i$ denotes the scale index) are processed via block causal attention, tokens within the same scale can attend to each other freely, while attention across scales follows a causal attention.
Figure 4: Multimodal versatile attention mechanism. $T$ denotes the text tokens. $U$ denotes the continuous visual tokens for multimodal understanding or reference image tokens for image editing. $G_{i}$ denotes the $i$-th scale discrete visual tokens for visual generation.
4 Model Training Pipeline
The training pipeline is divided into three stages and the instruction of the training recipe is shown in Tab. 1.
Figure 5: Overview of the training pipeline. In Stage 1, we first prepare a teacher model by training a two-layer MLP to connect InterViT [14] and the Qwen2.5 LLM [77]. This teacher model is then used to perform understanding distillation for the Und. FFN and the Patch Embedding layer. Simultaneously, we perform generation pretraining to optimize the Gen. FFN. All other parameters of the LLM remain frozen to preserve its pretrained language capabilities. In Stage 2 and 3, the entire model is unfrozen to conduct unified mid-training and supervised fine-tuning (SFT), respectively. The VAE component for visual generation is omitted from the figure for clarity.
4.1 Stage-1: Multilmodal Pretraining
The objective of this stage is to equip the OneCAT with foundational visual perception and generation abilities while maintaining the linguistic capabilities from the pretrained LLM. The core challenge is that, for visual perception, the Und. FFN is initialized from the weights of the LLM’s text-focused FFN. While this "warm start" benefits abstract reasoning, it inherently lacks pretrained visual knowledge, making the training process highly data-intensive. To address this limitation, we leverage the visual perception capabilities of an MLLM teacher and introduce an understanding distillation strategy to optimize Und. FFN that significantly enhances visual learning efficiency. In parallel, we conduct generation pretraining to optimize Gen. FFN.
4.1.1 Stage 1-1: Teacher Training
Rather than employing an off-the-shelf MLLM as teacher (e.g., Qwen2.5-VL [3]), we construct a custom teacher model to ensure parameter consistency between the LLM backbones of the teacher and student models, thereby improving distillation efficiency. Specifically, the teacher is built by connecting a pre-trained vision encoder (InterViT [14]) and a LLM (Qwen2.5 [77]) with a two-layer MLP as connector. During this sub-stage, we freeze the ViT and LLM, and exclusively train the MLP connector on a curated small-scale dataset of image-to-text caption pairs using the standard NTP loss.
4.1.2 Stage 1-2: Expert Pretraining
With the teacher model prepared, we proceed to train the OneCAT model. We keep the QKV Projection, Attention, and Text FFN frozen, and selectively optimize the task-specific modules: the Und. FFN and Patch Embedding Layer for multimodal understanding, and the Gen. FFN for text-to-image generation.
Hyperparameter / Config Stage 1-1 Stage 1-2 Stage 2 Stage 3
(Teacher Training) (Expert Pretraining) (Unified Mid-Training) (Unified SFT)
Learning Rate $2\times 10^{-3}$ $2\times 10^{-4}$ $2\times 10^{-5}$ $1\times 10^{-5}$
LR Scheduler Cosine Cosine Cosine Cosine
Weight Decay 0 0 0.01 0.01
Gradient Norm Clip 1.0 1.0 1.0 1.0
Batch Size 512 2048 512 256
Sequence Length 1024 1024 8192 16384
Number of Sample: Text-Only - - 40M 2M
Number of Sample: Und. 10M 436M 70M 11M
Number of Sample: Gen. - 52M 60M 3M
Number of Token (Total) 5B 0.3T 0.6T 57B
Token Ratio (T:U:G): 0:1:0 0:8:1 1:2:6 1:5:6
Resolution: Und. $\times448$ $\times448$ Native Native
Use thumbnail $\times$ $\times$ $\checkmark$ $\checkmark$
Resolution: Gen. - $\times256$
Dynamical
[…]
Dynamical
[…]
Number of Scales : Gen. - 7 $10\sim$13
Table 1: Detailed hyperparameter and configuration of the training recipe across different stages.
Understanding Distillation: We optimize the Und. FFN on a large-scale dataset of image-to-text caption pairs to acquire fundamental visual knowledge. Based on the specifically designed teacher, the training objective is a combination of the NTP loss ($\mathcal{L}_{\text{NTP}}$) and a distillation loss ($\mathcal{L}_{\text{Distill}}$). Specifically, $\mathcal{L}_{\text{NTP}}$ is the standard cross-entropy loss for autoregressive text generation. For distillation, instead of matching the final output logits, we align the student’s internal hidden states with those of the teacher model through deep feature-level matching. This strategy enables the student to not only mimic the teacher’s final prediction but also its intermediate computational patterns across all token positions (including both visual and text tokens) for better visual knowledge transfer. The distillation loss $\mathcal{L}_{\text{Distill}}$ is formulated as the sum of MSE losses between the hidden states of the student and teacher models over all $N$ transformer layers:
$\mathcal{L}_{\text{Distill}}=\sum_{n=1}^{N}\text{MSE}(\mathbf{h}_{S}^{(n)},\mathbf{h}_{T}^{(n)}),$ (1)
where $\mathbf{h}_{S}^{(n)}$ and $\mathbf{h}_{T}^{(n)}$ represent the hidden state outputs from the $n$-th transformer block of the student and teacher models, respectively. The final objective is thus:
$\mathcal{L}_{\text{Und}}=\mathcal{L}_{\text{NTP}}+\lambda\mathcal{L}_{\text{Distill}},$ (2)
where $\lambda=0.2$ is a balancing hyperparameter. Throughout distillation, all models process images at a fixed resolution of $448\times 448$ to balance computational load and the granularity of visual features.
Generation Pretraining: In parallel, we optimize the Gen. FFN on a delicate text-to-image generation dataset to enable the model to learn the spatial relationships and dependencies between multi-scale discrete visual tokens. We adopt the cross-entropy loss for next-scale prediction [63, 24] and the output image resolution is fixed to $\times256$ in this stage.
4.2 Stage-2: Unified Mid-Training
In the second stage, we unfreeze the entire model to achieve unified mid-training across multiple tasks (i.e., image-to-text, text-to-image, image editing, and text-only dialogues). At this stage, we incorporate the proposed scale-aware adapter (SAA), which is optimized with other modules of LLM together to extract scale-specific representation for enhanced image generation.
We introduce native resolution strategy for both understanding and generation in this stage. For multimodal understanding, the model is trained to process images at their original resolutions, thereby preserving fine-grained details and eliminating information loss. Additionally, a thumbnail of resolution $\times448$ is included to provide global visual context. For visual generation, the model is trained to generate images with different resolution and aspect ratios where the side lengths dynamically sampled from a range of 288 to 864 pixels, which significantly enhancing its generative versatility and real-world applicability.
4.3 Stage-3: Unified Supervised Fine-tuning
The final stage involves unified supervised fine-tuning (SFT) using a curated dateset of higher-quality data to enhance instruction-following and visual generation quality. The native resolution strategy was continued, with the size of generated image expanded to support side lengths between 288 to 1728 pixels, enabling higher-resolution results.
5 Data Setup
Stage-1: For the multimodal understanding, we curate a large-scale dataset of approximately 436 million image-text pairs, which is meticulously compiled and processed through comprehensive filtering and deduplication. This dataset is collected from two primary sources: (1) Public Available Image-Text Caption Pairs: We incorporate several publicly available, high-quality image-caption datasets, including Recap-DataComp-1B [38], Capsfusion [81], Detailed-Caption [39], SA1B-Dense-Caption [15], and Moondream2-COYO-5M-Captions [31]. (2) Re-captioned Image Datasets: We generate new image-text pairs by re-captioning large-scale public image collections. The source image datasets for this process include COYO700M [6], CC12M [7], CC3M [58], LAION-400M [57], and Zeor250M [74]. From this dataset, we randomly select a subset of 10 million samples to train the custom teacher.
For image generation, we construct a dataset of 52 million text-to-image samples after a rigorous filtering process to remove samples with low resolution or poor aesthetic scores. This collection consists of 1 million class-labeled images from ImageNet-1k [17], 20 million pairs from re-captioned public collections (i.e., COYO700M [6], LAION-400M [57] and CC12M [7]), and 30 million in-house synthetic images generated by FLUX. The overall training token ratio across multimodal understanding and visual generation samples in Stage-I is approximately 8:1.
Stage-2: In the unified mid-training, for multimodal understanding we leverage an in-house dataset of 70 million visual instruction samples. This dataset is specifically curated to be highly diverse tasks, including general VQA, detailed image captioning, OCR, multimodal reasoning(i.e., STEM problem-solving), knowledge, and visual grounding. For visual generation, we supplement the text-to-image samples of Stage-1 with a additional collection of 8 million image editing samples, resulting a total of 60 million visual generation samples. These additional image editing samples are sourced from several public image editing datasets, including AnyEdit [80], UltraEdit [88], HQ-Edit [30] and OmniEdit [68]. Additionally, we incorporate 40 million text-only instruction samples to preserve the language ability of LLM. To ensure a strong focus on visual generation in Stage-II, we oversample the visual generation data, resulting a final training token ratio of approximately 1 :2 :6 across text-only, multimodal understanding, and visual generation tasks, respectively.
OneCAT-1.5B OneCAT-3B
Base Model Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
Active Parameters 1.5B 3B
Total Parameters 4.5B 9B
Understanding Distillation
Teacher ViT InterViT-300M
Teacher LLM Qwen2.5-1.5B-instruct Qwen2.5-3B-instruct
Table 2: Model configurations for the two variants of OneCAT.
Stage-3: In the SFT stage, for multimodal understanding and text-only instruction, we construct a high-quality dataset of 13 million samples, combining 10 million from public MAmmoTH-VL dataset [22] with 3 million of our in-house synthetic instruction samples to further improve the reasoning abilities. For visual generation, we utilize a total of 3 million samples, aggregated from UniWorld [44], blip3o-60k [8], ShareGPT-4o-Image [10], and additional in-house synthetic data generated by GPT-4o and FLUX. The overall training token ratio across text-only, multimodal understanding, and visual generation is approximately 1 :5 :6.
6 Implementation details
Model Configurations: We conduct our experiments on two model variants, OneCAT-1.5B and OneCAT-3B. The OneCAT-1.5B model is based on Qwen2.5-1.5B-instruct [3] and contains 1.5B active parameters (4.5B total). The OneCAT-3B model is built upon Qwen2.5-3B-instruct [3] and utilizes 3B active parameters (9B total). For understanding distillation, the ViT of teacher model is InterViT-300M [14] for both two variants and the LLM of teacher model is Qwen2.5-1.5B-instruct and Qwen2.5-3B-instruct for OneCAT-1.5B and OneCAT-3B, respectively, to align with the LLM backbone of our OneCAT. Detailed instruction of the model configurations is shown in Tab. 2.
Table 3: Performance comparison across multiple multimodal understanding benchmarks. Higher scores are better, as indicated by the up-arrow ($\uparrow$). A-LLM denotes the number of activated LLM parameters, while Vis. indicates the parameter count of the vision encoder or tokenizer for multimodal understanding. Chameleon [62] does not report the parameter count of its vision tokenizer. slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding. Top-1 accuracy is reported (Best in bold, second best is underlined).
[…]
A-LLM Vis.
Encoder-based Understanding Only Models
InternVL2-2B [13] 1.8B 0.3B 73.4 76.2 58.9 86.9 - 74.1
InternVL2.5-2B [12] 1.8B 0.3B 74.3 79.2 60.9 88.7 - 74.9
Qwen2-VL-3B [3] 1.5B 0.6B 79.7 73.5 65.5 90.1 - 74.7
Qwen2.5-VL-3B [3] 3B 0.6B 79.3 84.0 77.1 93.9 - 81.6
Encoder-free Understanding Only Models
Mono-InternVL [49] 1.8B / 72.6 73.7 - - 59.5 68.6
EvE [18] 7B / 56.8 59.1 - - 62.6 61.0
EvEv2 [19] 7B / 71.1 73.9 - - 62.9 74.8
HoVLE [61] 2.6B / 70.9 78.6 55.7 86.1 64.9 73.0
Unified Models
Chameleon [62] 7B - 4.8 2.9 5.0 1.5 - 46.0
Emu3 [67] 8B 0.3B 64.7 68.6 43.8 76.3 60.3 70.0
Harmon-1.5B [71] 1.5B 0.9B - - - - 58.9 -
Show-o2-1.5B [76] 1.5B 0.5B - - - - 60.0 69.0
Janus-Pro-1.5B [11] 1.5B 0.3B - - - - 59.3 -
[…]
ILLUME+ [28] 3B 0.6B - 69.9 44.1 80.8 - 74.2
VILA-U [72] 7B 0.4B 60.8 - - - 60.8 -
Janus-Pro-7B [11] 7B 0.3B - - - - 62.0 -
Tar-7B [23] 7B 0.4B - - - - 61.3 -
Show-o2-7B [76] 7B 0.5B - - - - 63.1 78.6
[…]
Data Packing and Gradient Accumulation: To optimize workload balance across distributed processes and increase training throughput, we employ a data packing strategy that concatenates multiple variable-length samples into contiguous sequences. Furthermore, to manage the gradient contributions and token ratios between modalities as in Table 1, we utilize an uneven gradient accumulation strategy: prior to each optimizer step, we accumulate a distinct number of micro-batches’ gradients for the text and image generation tasks to obtain a gradient of desired token ratios. Such an approach provides fine-grained control over the effective batch sizes of different tasks, ensuring a balanced and stable joint-training.
Unbiased Global Batch Gradients: When training on $N$ distributed processes, naively averaging local loss can lead to biased gradients when per-process token counts vary. The ideal objective is to optimize the Global Batch Loss, defined as the loss summed over tokens for all micro-batches, normalized by the global token count, denoted as $T_{global}$. To this end, we first prefetch all micro-batches for the next optimizer step, enabling each process to compute the local token counts; a subsequent All-Reduce collective operation then aggregates these local token counts into the final global token count, i.e., $T_{global}$. Similar to [42], we then employ Global Batch Reduced Loss by dividing each micro-batch loss by the averaged token count per process, $\tfrac{T_{global}}{N}$, which can be shown that the final synchronized gradient for the subsequent optimizer step is mathematically equivalent to the gradient of the global batch loss, enabling training with unbiased gradients.
Table 4: Performance comparison across multiple multimodal understanding benchmarks. Higher scores are better, as indicated by the up-arrow ($\uparrow$). A-LLM denotes the number of activated LLM parameters, while Vis. indicates the parameter count of the vision encoder or tokenizer for multimodal understanding. Chameleon [62] does not report the parameter count of its vision tokenizer. slash (/) denotes that models do not require a vision encoder or tokenizer for multimodal understanding. Top-1 accuracy is reported (Best in bold, second best is underlined).
[…]
A-LLM Vis.
Encoder-based Understanding Only Models
InternVL2 [13] 1.8B 0.3B 1440 1877 73.2 34.3 44.6 46.4 71.6
InternVL2.5 [12] 1.8B 0.3B - 2138 74.7 43.6 60.8 51.3 -
Qwen2-VL [65] 1.5B 0.6B - 1872 74.9 41.1 49.5 43.0 -
Qwen2.5-VL [3] 3B 0.6B - 2157 79.1 53.1 61.8 62.3 -
Encoder-free Understanding Only Models
Mono-InternVL [49] 1.8B / - 1875 65.5 33.7 40.1 45.7 67.4
EvE [18] 7B / - 1628 52.3 32.6 25.7 - 64.6
EvEv2.0 [19] 7B / - 1709 66.3 39.3 45.0 - 71.4
HoVLE [61] 2.6B / - 1864 71.9 33.7 44.3 46.2 70.7
Unified Models
Chameleon [62] 7B - - - 35.7 28.4 8.3 - 30.6
Emu3 [67] 8B 0.3B - - 58.5 31.6 37.2 - 68.2
Harmon [71] 1.5B 0.9B 1155 1476 65.5 38.9 - - 67.1
Show-o2 [76] 1.5B 0.5B 1450 - 67.4 37.1 - - 65.6
Janus-Pro [11] 1.5B 0.3B 1444 - 75.5 36.3 39.8 - -
[…]
ILLUME+ [28] 3B 0.6B $14$ - 80.8 44.3 40.3 - 73.3
VILA-U [72] 7B 0.4B 1401 - - - 33.5 - 59.0
Janus-Pro [11] 7B 0.3B 1567 - 79.2 41.0 50.0 - -
Tar [23] 7B 0.4B 1571 1926 74.4 39.0 - - -
Show-o2 [76] 7B 0.5B 1620 - 79.3 48.9 - - 69.8
[…]
Figure 6: Text-to-Image comparison.
Figure 7: Image-Editing comparison.
7 Evaluation
7.1 Multimodal understanding.
We evaluate our model on 12 public Multimodal understanding benchmarks spanning diverse capabilities: MMbench [47], MME [79], MMMU [83], MM-Vet [82], and SEED [35] assess general multimodal perception and reasoning. MathVista [48] focuses on mathematical reasoning. TextVQA [60], ChartQA [50], InfoVQA [52], and DocVQA [51] evaluate OCR and text-related visual question answering; while GQA [29] evaluates visual scene understanding and AI2D [33] evaluates scientific diagram comprehension.
Table 5: Performance comparison on the GenEval [21] benchmark. The dagger () indicates methods that employ an LLM for prompt rewriting. Best in bold, second best is underlined.
[…]
Generation-only Models
SDXL [54] 0.98 0.74 0.39 0.85 0.15 0.23 0.55
DALL-E 3 [59] 0.96 0.87 0.47 0.83 0.43 0.45 0.67
Infinity [24] - 0.85 - - 0.49 0.57 0.73
SD3-Medium [20] 0.99 0.94 0.72 0.89 0.33 0.60 0.74
FLUX.1-dev [34] 0.98 0.93 0.75 0.93 0.68 0.65 0.82
Unified Models
Chameleon-7B [62] - - - - - - 0.39
Transfusion-7B [90] - - - - - - 0.63
Emu3-8B [67] 0.99 0.81 0.42 0.80 0.49 0.45 0.66
ILLUME+ 3B [28] 0.99 0.88 0.62 0.84 0.42 0.53 0.72
Harmon-1.5B [71] 0.99 0.86 0.66 0.85 0.74 0.48 0.76
Show-o2-7B [76] 1.00 0.87 0.58 0.92 0.52 0.62 0.76
Janus-Pro-7B [11] 0.99 0.89 0.59 0.90 0.79 0.66 0.80
Mogao-7B [41] 1.00 0.97 0.83 0.93 0.84 0.80 0.89
BLIP3-o-8B [8] - - - - - - 0.84
Tar-7B [23] 0.99 0.92 0.83 0.85 0.80 0.65 0.84
UniWorld-V1-20B [43] 0.99 0.93 0.79 0.89 0.49 0.70 0.80
UniWorld-V1-20B [43] 0.98 0.93 0.81 0.89 0.74 0.71 0.84
BAGEL-7B [16] 0.99 0.94 0.81 0.88 0.64 0.63 0.82
BAGEL-7B [16] 0.98 0.95 0.84 0.95 0.78 0.77 0.88
[…]
As shown in Tab. 3 and Tab. 4, we compare OneCAT with three types of models: encoder-based understanding-only models, encoder-free understanding-only models, and unified MLLMs. Our OneCAT-3B model demonstrates superior performance, significantly outperforming all existing encoder-free understanding-only MLLMs, e.g., HoVLE [61] and EvEv2 [19], across nearly all benchmarks. For instance, on OCR-related tasks including TextVQA (73.9), ChartQA (81.2), InfoVQA (64.8), and DocVQA (91.2), OneCAT-3B achieves new state-of-the-art results among encoder-free models. It also excels in general vision-language benchmarks such as MME-P (1630), MMBench (78.8), and MM-Vet (52.2).
Moreover, OneCAT-3B outperforms recent unified MLLMs that rely on external vision encoders or tokenizers—such as Janus-Pro-7B [11] (using SigLIP [84]) and Tar-7B [23] (using SigLip2 [64])—despite activating fewer parameters. Compared to top-tier encoder-based understanding-only models like Qwen2.5-VL-3B [3], our model exhibits a slight performance gap, which we primarily attribute to differences in the scale and quality of training data. Specifically, Qwen2.5-VL was trained on 4T tokens, whereas our OneCAT was trained on only 0.5T tokens for multimodal understanding. We believe this gap can be bridged in the future by scaling up the pretraining data and incorporating more higher-quality instruction data.
7.2 Visual Generation.
We evaluate our model on three widely-used visual generation benchmarks: two for text-to-image generation, GenEval [21] and DPG-Bench [27], and one for instruction-based image editing, ImgEdit [78]. We follow previous works [10, 41, 16] to use Classifier-free guidance(CFG) [25] to enhance visual generation quality. During training, we randomly drop tokens of conditional text and reference image with probabilities 0.1, 0.1, respectively. During inference, we combines conditional and unconditional predicted logits to produce outputs that better adhere to the given conditions. To ensure a fair comparison, we adhere strictly to the original raw prompts for the GenEval benchmark, unlike some previous works that employ LLM-based prompt rewriting to enhance performance. As shown in Tab.5, 6, and 7, our OneCAT-3B model demonstrates highly competitive performance across all tasks.
On GenEval (Tab. 5), which evaluates fine-grained instruction following over object counts, colors, and spatial relationships, OneCAT-3B achieves a SOTA overall score of 0.90. This performance surpasses most all unified models, including the strong baseline BAGEL-7B (0.88 with prompt rewriting) and Mogao-7B (0.89 with prompt rewriting). Notably, OneCAT-3B excels in challenging categories such as Position and Color Attribute, where it achieves the best performance (0.84 and 0.80 respectively), showcasing its superior ability to interpret complex spatial and attribute-based instructions.
On DPG-Bench (Tab. 6), a benchmark focused on compositional text-to-image generation, OneCAT-3B attains a strong overall score of 84.53. This result is highly competitive among unified models, outperforming strong counterparts like Janus-Pro-7B (84.19) and Mogao-7B (84.33).
On ImgEdit-Bench (Tab. 7), a challenging and diverse image editing benchmark, OneCAT-3B achieves an overall score of 3.43. This places it firmly among the top-performing unified models and significantly outperforms many specialized editing models. OneCAT-3B demonstrates exceptional capabilities in categories requiring precise local and global adjustments, securing the top scores in Adjust (3.70), Extract (2.42), and Background (3.79) manipulation. This highlights the effectiveness of our model’s ability to condition its generation on fine-grained visual cues from a reference image.
Figures 6 and 7 present qualitative comparisons for the text-to-image and image-editing tasks, respectively. OneCAT-3B exhibits leading instruction-following and world-understanding capabilities among open-source models. For example, in the fourth row of Figure 6, it is the only open-source model that correctly generates exactly four characters in a pixel-art style; in contrast, BAGEL-7B produces five characters, and Janus-Pro-7B fails to render pixel art. Similarly, in the last row of Figure 7, only OneCAT-3B produces an image with the correct lighting conditions.
In summary, the strong instruction-following capability in both text-to-image generation and image editing tasks highlights the effectiveness and versatility of our architectural design.
Table 6: Performance comparison on the DPG-Bench [27] benchmark. Best in bold, second best is underlined.
[…]
Generation-only Models
Hunyuan-DiT [40] 84.59 80.59 88.01 74.36 86.41 78.87
Playground v2.5 [36] 83.06 82.59 81.20 84.08 83.50 75.47
PixArt-$\Sigma$ [9] 86.89 82.89 88.94 86.59 87.68 80.54
DALL-E 3 [59] 90.97 89.61 88.39 90.58 89.83 83.50
Infinity [24] 93.11 - - 90.76 - 83.46
SD3-Medium [20] 87.90 91.01 88.83 80.70 88.68 84.08
FLUX.1-dev [34] 82.10 89.50 88.80 91.10 89.40 84.00
Unified Models
Emu3-8B [67] - - - - - 81.60
Janus-Pro-7B [11] 86.90 88.90 89.40 89.32 89.48 84.19
Mogao-7B [41] 82.37 90.03 88.26 93.18 85.40 84.33
BLIP3-o-8B [8] - - - - - 81.60
Tar-7B [23] 83.98 88.62 88.05 93.98 84.86 84.19
Show-o2-7B [76] 89.00 91.78 89.96 91.81 91.64 86.14
[…]
Table 7: Comprehensive comparison on ImgEdit-Bench [78] showing performance across nine editing categories. Higher scores are better for all metrics. Best in bold, second best is underlined.
Model Add Adjust Extract Replace Remove Background Style Hybrid Action Overall
Editing-only Models
MagicBrush [85] 2.84 1.58 1.51 1.97 1.58 1.75 2.38 1.62 1.22 1.90
Instruct-Pix2Pix [5] 2.45 1.83 1.44 2.01 1.50 1.44 3.55 1.20 1.46 1.88
AnyEdit [32] 3.18 2.95 1.88 2.47 2.23 2.24 2.85 1.56 2.65 2.45
UltraEdit [87] 3.44 2.81 2.13 2.96 1.45 2.83 3.76 1.91 2.98 2.70
Step1X-Edit [46] 3.88 3.14 1.76 3.40 2.41 3.16 4.63 2.64 2.52 3.06
ICEdit [86] 3.58 3.39 1.73 3.15 2.93 3.08 3.84 2.04 3.68 3.05
Unified Models
OmniGen [73] 3.47 3.04 1.71 2.94 2.43 3.21 4.19 2.24 3.38 2.96
OmniGen2 [70] 3.57 3.06 1.77 3.74 3.20 3.57 4.81 2.52 4.68 3.44
BAGEL-7B [16] 3.56 3.31 1.70 3.30 2.62 3.24 4.49 2.38 4.17 3.20
UniWorld-V1-20B [43] 3.82 3.64 2.27 3.47 3.24 2.99 4.21 2.96 2.74 3.26
OneCAT-3B 3.65 3.70 2.42 3.92 3.00 3.79 4.61 2.23 3.53 3.43
7.3 Comparison of Inference Efficiency
In addition to its strong performance, our model’s architectural design also yields significant improvements in inference efficiency for handling both high-resolution image input and output. We evaluated the efficiency of OneCAT-3B in two distinct phases: prefilling and generation.
In the prefilling phase, we measure the Time to the First Token (TTFT) to assess the computational cost of processing the input. As shown in Tab. 8, thanks to our pure decoder-only architecture that eliminates the need for a separate ViT encoder, OneCAT-3B demonstrates a substantial efficiency advantage, particularly when handling high-resolution images. Compared to QwenVL-3B, as the image resolution increases from $768\times 768$ to $1792\times 1792$, the reduction in TTFT for OneCAT-3B grows sharply from 50.4% to 61.4%. This provides strong evidence for the effectiveness of our design, especially for high-resolution inputs.
In the generation phase, we evaluated the total inference time for text-to-image (T2I) generation and image editing. As detailed in Tab. 9, OneCAT-3B’s inference speed far surpasses that of the strong diffusion-based baseline, BAGEL-7B. For instance, when generating a $1024\times 1024$ resolution image, OneCAT-3B’s T2I and editing inference times are merely 2.85s and 4.61s, respectively, making it approximately $10\times$ faster than BAGEL-7B. This significant advantage stems from our pioneering multi-scale autoregressive generation mechanism inside LLM that enables parallel token generation, which drastically reduces the number of required decoding steps compared to iterative diffusion models.
In summary, OneCAT achieves exceptional inference efficiency in both the prefilling and generation stages. This highlights the immense potential and practical value of our unified autoregressive framework for building efficient and powerful large multimodal models.
Table 8: Efficiency comparison of OneCAT-3B and QwenVL-3B. Models are tested based on one NVIDIA H800 GPU. We report the time to the first token (TTFT) to measure the reduction of computational cost of prefilling phase. 256* denotes the number of visual tokens for thumbnail.
Model Resolution of Input Image #Input Text Tokens #Input Visual Tokens TTFT(s) Reduction
Qwen2.5-VL-3B $768\times 768$ 24 731 0.135 50.4%
OneCAT-3B $768\times 768$ 24 731 $256^{*}$ 0.067
Qwen2.5-VL-3B $1024\times 1024$ 24 1395 0.216 57.4%
OneCAT-3B $1024\times 1024$ 24 1395 $256^{*}$ 0.092
Qwen2.5-VL-3B $1792\times 1792$ 24 4098 0.583 61.4%
OneCAT-3B $1792\times 1792$ 24 4098 $256^{*}$ 0.225
Table 9: Generation efficiency comparison of OneCAT-3B and BAGEL. Models are tested based on one NVIDIA H800 GPU. We report total inference time to measure the Text-to-Image(T2I) and Image-Editing efficiency.
Model Resolution of Generated Image T2I Infer. Time (s) Edit Infer. Time (s)
BAGEL-7B $512\times 512$ 8.76 13.45
OneCAT-3B $512\times 512$ 1.40 2.03
BAGEL-7B $1024\times 1024$ 26.29 46.44
OneCAT-3B $1024\times 1024$ 2.85 4.61
8 Conclusion
In this work, we presented OneCAT, a pure decoder-only unified multimodal model that seamlessly integrates understanding, generation, and editing within a single, streamlined architecture. By eliminating external encoders and tokenizers, employing a modality-specific MoE design, and introducing a multi-scale autoregressive generation mechanism, OneCAT achieves strong performance across a wide range of benchmarks while significantly improving inference efficiency. Our results demonstrate the viability and advantages of a first-principles approach to multimodal modeling, offering a powerful new baseline for future research and applications in general-purpose multimodal intelligence.